Prove that cos π/5.cos 2π/5=1/4.
And cos π/5 -cos 2π/5=1/2
 
 

Grade:11

1 Answers

anmol
5 years ago
quite easy,
1/2(2cospi/5cos2pi/5)-1/4
2cosAcosB=cos(A+B)+cos(A-B)
cos3pi/2+cospi/5-1/2
cos(pi-2pi/5)+cospi/5=1/2
cospi/5-cos2pi/5=1/2,